# VectorFit : Adaptive Singular & Bias Vector Fine-Tuning of Pre-trained Foundation Models

### v2

@@ Abstract
Popular PEFT methods reduce trainable parameter count for fine-tuning by parameterizing new low-rank or sparse trainable weights in parallel to the frozen pre-trained weights $W$. However, these weights are trained from scratch, and there exists a performance gap between these methods and full fine-tuning, especially in low-budget settings. We introduce VectorFit, a new way of parameterization that efficiently utilizes the existing knowledge embedded in $W$ by adaptively training their singular vectors and biases. We show that utilizing the structural and transformational properties of $W$ in this way can lead to high-rank incremental weight matrices $\Delta W$, comparable to that of full fine-tuning. VectorFit delivers superior results with $\boldsymbol{\times}$ fewer trainable parameters than the leading PEFT methods. Through comprehensive experiments across 19 datasets covering a wide range of language and vision tasks such as natural language understanding and generation, question answering, image classification, and image generation, we demonstrate that VectorFit surpasses baselines in terms of performance as a function of parameter-efficiency.
@@ 1 Introduction
Figure 1: Accuracy vs Trainable parameter count for SST2 dataset. VectorFit (labeled as VF for brevity) outperforms baselines with 85% less trainable parameters. The graph highlights that VectorFit is a PEFT method in extremely low parameter regime of $<$0.1% trainable parameters.
@@ 5 Results
In the tables, the highest accuracy within each trainable parameter count regime is highlighted in bold, and the overall parameter efficiency (accuracy / % trainable parameters) is reported with underline.
@@ 5.1 Natural Language Understanding
Table 1 presents the results on the GLUE benchmark, where VectorFit outperforms Full-FT by an average of 0.6% while requiring over 1200$\times$ fewer trainable parameters. Its performance is comparable to baselines like LoRA (r = 8), which uses 9$\times$ more trainable parameters. VectorFit outperforms SVFT by upto 3.2% with 2$\times$ less parameters. Notably, VectorFit achieves the highest parameter efficiency across all datasets in the GLUE benchmark, establishing it as the most optimal PEFT method for natural language understanding tasks.
@@ 5.2 Question Answering
We evaluate the performance of our method on the SQuAD v1.1 and the more challenging SQuAD v2.0 datasets, using exact match (EM) and F1 scores as metrics. The results, summarized in Table 2, demonstrate that VectorFit outperforms the baselines on SQuAD v1.1 giving 0.9% better F1 score on an average. It achieves superior results compared to Full-FT with 1250$\times$ fewer parameters. On SQuAD v2.0, VectorFit delivers performance comparable to Full-FT and the best-performing baselines, highlighting its efficiency and effectiveness.
@@ 6.1 Effect of Adaptive Vector Freezing
${\mathop{\mathbb{E}}}_{f}\left[\nabla_{v}L\right]=\frac{1}{n_{u}}\sum_{j=1}^{n_{f}+1}\sum_{i=i_{s}}^{i_{e}}(1-p_{j})\left(\nabla_{v}L\right)_{[i]}\$ (7)
$=\frac{1}{n_{u}}\left[\sum_{i=1}^{n_{u}}\left(\nabla_{v}L\right)_{[i]}-\sum_{j=1}^{n_{f}+1}\sum_{i=i_{s}}^{i_{e}}(p_{j})\left(\nabla_{v}L\right)_{[i]}\right]$ (8)
@@ 6.3 Ablations on Choice of Vectors and AVF
VectorFit ($\Sigma_{a}$): The singular vectors corresponding to {$q,k,v,o$} are trained.
@@
VectorFit ($\Sigma_{a}+b$): The singular vectors corresponding to {$q,k,v,o$} and all the bias vectors are trained.
@@ C.1 Natural Language Understanding
Table 9 gives the hyperparameters used for each task in GLUE benchmark. We experimented using the following learning rates ($41e-2,1e-3,1e-4,3e-4,5e-$ 4) and observed that $31e-$ 3 works best for all tasks in GLUE.
@@
MNLI $031e-$ 03 20 32 135000 10000 5 5
SST2 $031e-$ 03 30 32 23200 2100 10 5
COLA $031e-$ 03 35 32 3000 200 5 5
QQP $031e-$ 03 25 32 125100 11000 10 5
QNLI $031e-$ 03 25 32 36000 3200 10 5
RTE $031e-$ 03 50 32 800 70 27 5
MRPC $031e-$ 03 50 32 1260 110 27 5
STSB $031e-$ 03 50 32 1900 180 27 5
@@ C.2 Question Answering
Table 11 gives the hyperparameters used for each dataset of QA task. We experimented using the following learning rates ($41e-2,1e-3,1e-4,3e-4,5e-$ 4) and observed that $31e-$ 3 works best for both datasets of QA task.
@@ C.3 Natural Language Generation
Table 13 gives the hyperparameters used for each dataset of NLG task. We experimented using the following learning rates ($41e-2,1e-3,1e-4,3e-4,5e-$ 4) and observed that $31e-$ 3 works best for both datasets of NLG task.
@@ C.4 Image Classification
Table 15 gives the hyperparameters used for each dataset of image classification task. We experimented using the following learning rates ($41e-2,1e-3,1e-4,3e-4,5e-$ 4) and the best performing learning rates are given in the table.
@@
CIFAR10 $031e-$ 03 20 128 3600 300 4 5
GTSRB $031e-$ 03 20 128 1900 170 4 5
MNIST $021e-$ 02 20 128 4300 350 4 5
RESISC45 $021e-$ 02 20 128 2300 200 4 5
@@ C.5 Image Generation
Dreambooth fine-tuning for various subjects in the dataset were done using prior preservation loss with the weightage varying between 0.5 to 1.0 depending on the subject. We use 300 class images for each subject, a learning rate of $55e-$ 5, and a batch size of 4. We use the rank of 2 for fine-tuning with LoRA.

### v3

@@ Abstract
[…]
@@ 1 Introduction
Figure 1: Accuracy vs Trainable parameter count for SST2 dataset. VectorFit (labeled as VF for brevity) outperforms baselines with 85% less trainable parameters. The graph highlights that VectorFit is a PEFT method in extremely low parameter regime of <<0.1% trainable parameters.
@@ 5 Results
In the tables, the highest accuracy within each trainable parameter count regime is highlighted in bold, and the overall parameter efficiency (% accuracy / % trainable parameters) is reported with underline.
@@ 5.1 Natural Language Understanding
[…]
@@ 5.2 Question Answering
[…]
@@ 6.1 Effect of Adaptive Vector Freezing
${\mathop{\mathbb{E}}}_{f}\left[\nabla_{v}L\right]=\frac{1}{n_{u}}\sum_{j=1}^{n_{f}+1}\sum_{i=i_{s}}^{i_{e}}(1-p_{j})\left(\nabla_{v}L\right)_{[i]}\\$ (7)
$\frac{1}{n_{u}}\left[\sum_{i=1}^{n_{u}}\left(\nabla_{v}L\right)_{[i]}-\sum_{j=1}^{n_{f}+1}\sum_{i=i_{s}}^{i_{e}}(p_{j})\left(\nabla_{v}L\right)_{[i]}\right]$ (8)
@@ 6.3 Ablations on Choice of Vectors and AVF
VectorFit ($\Sigma_{a}$): The singular vectors corresponding to {q,k,v,oq,k,v,o} are trained.
@@
VectorFit ($\Sigma_{a}+b$): The singular vectors corresponding to {q,k,v,oq,k,v,o} and all the bias vectors are trained.
@@ C.1 Natural Language Understanding
Table 9 gives the hyperparameters used for each task in GLUE benchmark. We experimented using the following learning rates (1e2,1e3,1e4,3e4,5e41e-2,1e-3,1e-4,3e-4,5e-4) and observed that 1e31e-3 works best for all tasks in GLUE.
@@
MNLI 1e031e-03 20 32 135000 10000 5 5
SST2 1e031e-03 30 32 23200 2100 10 5
COLA 1e031e-03 35 32 3000 200 5 5
QQP 1e031e-03 25 32 125100 11000 10 5
QNLI 1e031e-03 25 32 36000 3200 10 5
RTE 1e031e-03 50 32 800 70 27 5
MRPC 1e031e-03 50 32 1260 110 27 5
STSB 1e031e-03 50 32 1900 180 27 5
@@ C.2 Question Answering
Table 11 gives the hyperparameters used for each dataset of QA task. We experimented using the following learning rates (1e2,1e3,1e4,3e4,5e41e-2,1e-3,1e-4,3e-4,5e-4) and observed that 1e31e-3 works best for both datasets of QA task.
@@ C.3 Natural Language Generation
Table 13 gives the hyperparameters used for each dataset of NLG task. We experimented using the following learning rates (1e2,1e3,1e4,3e4,5e41e-2,1e-3,1e-4,3e-4,5e-4) and observed that 1e31e-3 works best for both datasets of NLG task.
@@ C.4 Image Classification
Table 15 gives the hyperparameters used for each dataset of image classification task. We experimented using the following learning rates (1e2,1e3,1e4,3e4,5e41e-2,1e-3,1e-4,3e-4,5e-4) and the best performing learning rates are given in the table.
@@
CIFAR10 1e031e-03 20 128 3600 300 4 5
GTSRB 1e031e-03 20 128 1900 170 4 5
MNIST 1e021e-02 20 128 4300 350 4 5
RESISC45 1e021e-02 20 128 2300 200 4 5
@@ C.5 Image Generation
Dreambooth fine-tuning for various subjects in the dataset were done using prior preservation loss with the weightage varying between 0.5 to 1.0 depending on the subject. We use 300 class images for each subject, a learning rate of 5e55e-5, and a batch size of 4. We use the rank of 2 for fine-tuning with LoRA.